# Article Category

### v2

@@ 3.3 ZONOTOPES
Translating a zonotope is equivalent to translating the centre. Linearly mapping by $M\in\mathbb{R}^{{\color[rgb]{0,0,0}m}\times d}$ yields [46, Eq. 7b]
@@
A support function of a set describes the farthest extent of the set in a given direction. The support function of a zonotope in direction $v\in\mathbb{R}^{d}$ is [61, Lemma 1]
@@
A point $p\in\mathbb{R}^{d}$ is contained in a zonotope if [62, Eq. 6]
@@
Determining the containment of a zonotope in another zonotope is co-NP complete [62], but a sufficient condition for $\mathcal{Z}_{1}\subseteq\mathcal{Z}_{2}$ is [63, Eq. 15]
1 1$\geq$1 ≥ $\!\min_{\gamma\in\mathbb{R}^{n_{2}},\Gamma\in\mathbb{R}^{n_{2}\times n_{1}}}$ $\left\lVert\begin{bmatrix}\Gamma&\gamma\end{bmatrix}\right\rVert_{\infty}$ (8a)
@@ 4 PROBLEM STATEMENT
Our work considers constrained Markov decision processes $(\mathcal{S},\mathcal{A},P_{f},r,\mathcal{A}_{s})$ of the following elements:
  • a feasible state set $\mathcal{S}\subset\mathbb{R}^{d_{\mathcal{S}}}$,
  • a feasible action set $\mathcal{A}\subset\mathbb{R}^{d}$,
@@
We seek a safeguarded, stochastic policy that maximises the expected, discounted return over a finite horizon $N$
$\pi^{*}(a|s)=\!\operatorname*{argmax}_{\pi(a|s)}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a_{i}\sim\pi(a_{i}|s_{i})\\ s_{i+1}\sim P_{f}\end{subarray}}\sum_{i=0}^{N}\delta^{i}\,r(s_{i},{a_{s}}_{i},s_{i+1})$ (9)
with the safe action ${a_{s}}_{i}{\color[rgb]{0,0,0}=g_{\mathcal{A}_{s}}(a_{i})}$, the continuously differentiable safeguard $g_{\mathcal{A}_{s}}\colon\mathcal{A}\to\mathcal{A}_{s}$, and discount factor $\delta\in(0,1{\color[rgb]{0,0,0}]}$.
4.1 ENSURING COMPUTATIONAL TRACTABILITY
@@
Safeguarding is computationally cheap when the problem setting explicitly provides such a representation for the safe action set. However, safeguarding might be computationally intractable if the safe action set needs to be constructed from a safe state set, as specified in Equation˜2, even if the safe state set is available as a closed, convex set. In these cases, we assume that the next state set can be derived using only computations adhering to disciplined convex programming [64]. This allows to replace any constraint on a safe action $a_{s_{i}}\in\mathcal{A}_{s}$ by the state constraint $\mathcal{S}_{i+1}(a_{s_{i}},s_{i})\subseteq\mathcal{S}_{s}$.
@@
One method we discuss in particular is obtaining safe state zonotopes via robust control invariant sets [46], which guarantee the existence of an invariance-enforcing controller that can keep all future states within the safe set. This is achieved by enclosing the transition distribution at the current state by a linear transition function with a noise zonotope $\mathcal{W}=\left<c_{\mathcal{W}},G_{\mathcal{W}}\right>\subset\mathbb{R}^{d_{\mathcal{S}}}$. Such an enclosure can, for example, be obtained using reachset-conformant identification [44], which bounds the remainder of the Taylor expansion. Such a transition function allows a linear computation of the next state set
$\mathcal{S}_{i+1}(a_{i},s_{i})=Ma_{i}+\left<c+c_{\mathcal{W}},G_{\mathcal{W}}\right>\,,$ (10)
where $c$ is the offset an $M$ the Jacobian of the linearisation.
@@ 5 METHOD
Figure˜1 shows the general framework for provably safe, analytic gradient-based reinforcement learning. For any policy output, we apply safeguards that map the unsafe action $a_{i}$ to the safe action ${a_{s}}_{i}$. The safe action is executed in the environment, yielding the next state $s_{i+1}$ and reward $r_{i}$. To train the policy, we calculate the gradient of the reward with respect to the policy output
$\frac{\partial r_{i}}{\partial a_{i}}=\left(\frac{\partial r_{i}}{\partial{a_{s}}_{i}}+\frac{\partial r_{i}}{\partial s_{i+1}}\frac{\partial s_{i+1}}{\partial{a_{s}}_{i}}\right)\frac{\partial{a_{s}}_{i}}{\partial a_{i}}\,.$ (11)
@@
Figure 1: The forward (top) pass of the provably safe policy optimisation from analytic gradients describes the integration of the safeguard in-between the policy and environment. The backward pass (bottom) visualises, how we utilise backpropagation to differentiate through the environment and safeguard to obtain the reward gradient with respect to the policy action. It also highlights how this process in principle requires the unrolling of the previous trajectory.
The following subsections detail the safeguard. First, we formulate generally required and desirable properties in the aforementioned differentiable setting. Then, we introduce the two safeguards used in this work: boundary projection (BP) [20] and ray mask (RM) [12].We structure their introduction by first explaining the principle of the safeguard, then analysing its properties, and finally present our modifications.
5.1 REQUIRED AND DESIRED PROPERTIES
A safeguard for our setting must be provably safe as described in Section˜3.2. To allow for backpropagation, it must also guarantee the existence of a Clark generalised derivative [69] everywhere, which implies that the safeguarding is at least of class $C^{0}$.
Beyond the required, there are additional desired properties. First, the safeguarding should be of class $C^{1}$ and provide full-rank Jacobians $\frac{\partial a_{s}}{\partial a}$ everywhere. A rank-deficient Jacobian can diminish the learning signal by reducing its effective dimensionality and incurring information loss. Second, the number of interventions by the safeguarding should be minimal throughout training and inference. Minimal interventions also serve the last desired property of fast computation. In summary, the safeguard must
@@ 5.1 REQUIRED AND DESIRED PROPERTIES
    be of class $C^{1}$ and provide full-rank Jacobians everywhere, i.e. be a local diffeomorphism,
@@
In the following, we present two safeguards that offer different trade-offs between the desired properties. We summarise the properties of the safeguards in Table˜1.
@@
For $\mathcal{A}_{s}$ 1 Quadratic Problem 1 Linear Problem
For $\mathcal{S}_{s}$wedge $P_{f,lin}$ 1 Quadratic Problem (1 Conic $\vee$ 1 Quadratic) $\wedge$ 1 Linear Problem
5.2 BOUNDARY PROJECTION
The boundary projection safeguard, which was proposed in [20], maps any action to the closest safe action. By definition, it therefore only affects unsafe actions, which are mapped to the nearest boundary point in the safe action set. In Euclidean space, this corresponds to an orthogonal projection to the boundary of the safe action set. We show an exemplary mapping with boundary projection from an unsafe action $a$ to a safe action $a_{s}$ in Figure˜2.
@@ 5.2 BOUNDARY PROJECTION
The safeguard $g_{\mathcal{A}_{s},\text{BP}}(a)$ provides the safe action by solving
@@
5.2.1 PROPERTIES
The optimisation problem is always solvable given Equation˜1, and the constraint Equation˜12b satisfies Property˜P1. The implicit function theorem [49] provides the Jacobian of the solution mapping to fulfil Property˜P2.
The distance between the initial and mapped action decreases smoothly with the distance from the unsafe action to the boundary until it is zero for safe actions. However, the mapping location can change abruptly between unsafe actions on different sides of the edges of the safe set. This leads to a jump in the gradient landscape, such that it is only $C^{1}$ almost everywhere. In addition, all unsafe actions along a normal vector to the safe action set are projected to the same safe action on the boundary. Formally, any safe action on the boundary $a_{s,\partial\mathcal{A}_{s}}\in\partial\mathcal{A}_{s}$ is the solution to Equation˜12 for all actions
$a=a_{s,\partial\mathcal{A}_{s}}+t\cdot v\quad\forall t,\,\forall v$ (13)
with the normal, unit vectors $v\in\{v\mid v\cdot(a_{s}-a_{s,\partial\mathcal{A}_{s}})\leq 0,\forall a_{s}\in\mathcal{A}_{s}\}\setminus\{0\}$ and a positive scalar $t$. Therefore, the safeguard cannot propagate gradient components in the mapping direction, such that
$\left(\frac{\partial r}{\partial a_{s}}\frac{\partial a_{s}}{\partial a}\right)^{T}v=0$ (14)
which is especially problematic for gradients parallel to $v$, that occur when the optimal action is safe. In such a case, boundary projection eliminates the gradient, keeping the optimisation stuck indefinitely.
Lemma 1.
Let $\mathcal{A}_{s}$ be the known zonotope $c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\right>$ with generator matrix $G_{\mathcal{A}_{s}}\in\mathbb{R}^{d\times n}$, where ($n\geq d$), such that strong duality holds for Equation˜12. Then, for safe actions, the Jacobian of Equation˜12 is the identity and thus has full rank. The Jacobian is a projection matrix of rank at most $d-1$ for unsafe actions.
We obtain a proof by differentiating through the KKT conditions, which we provide in Section˜7.6. Consequently, boundary projection does not fulfil Property˜P3.
Boundary projection only intervenes for the required unsafe actions and therefore adheres to Property˜P4. If the safe action set is explicitly available as a zonotope, Equation˜7 provides the containment constraint Equation˜12b. For a linearised transition distribution the constraint becomes $\mathcal{S}_{i+1}(a_{i},s_{i})\subseteq\mathcal{S}_{s}$, which is Equation˜8. Both yield quadratic problems, which compute quickly; see Property˜P5.
5.2.2 MODIFICATIONS
To regain a gradient component in the mapping direction and improve Property˜P3, we augment the policy loss function $l_{r}(a_{s},s)$ with a regularisation term [18, Eq. 16]
@@ 5.2.2 MODIFICATIONS
The corresponding gradient
$\frac{\partial l}{\partial a}=\frac{\partial l_{r}}{\partial a}+2c_{d}(a_{s}-a)(\frac{\partial a_{s}}{\partial a}-I)$ (16)
points along the projection direction. The coefficient $c_{d}$ scales the regularisation to remain small relative to the original loss $l_{r}(a_{s},s)$ yet large enough to produce a meaningful gradient component in the mapping direction. If $c_{d}$ is too large, the gradient may become uninformative or vanish for unsafe actions. Vanishing may occur if a hyperbolic tangent function squashes the policy output to adhere to a normalised action space. An excessive $c_{d}$ can yield large upstream gradients, which saturate the hyperbolic tangent. In addition to gradient augmentation, the regularisation encourages the policy to favour safe actions from the start, which is desirable for Property˜P4.
5.3 RAY MASK
Figure 3: The ray mask maps actions towards the safe centre in proportion to the safety domain length , the feasible domain length, and the distance from the action to the safe centre.
The ray mask technique [12] maps every action radially towards the centre of the safe action set $c_{\mathcal{A}_{s}}$, as shown in Figure˜3. Using the unit vector $d_{a}=\frac{a-c_{\mathcal{A}_{s}}}{\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert}$, we define the distances from the safe action set centre to the initial action and the boundaries of the safe and feasible action sets as
$\lambda_{a}$ $=\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert$ (17)
$\lambda_{\mathcal{A}_{s}}$ $=\sup\{\lambda\geq 0:c_{\mathcal{A}_{s}}+\lambda d\in\mathcal{A}_{s}\}$ (18)
$\lambda_{\mathcal{A}}$ $=\sup\{\lambda\geq 0:c_{\mathcal{A}_{s}}+\lambda d\in\mathcal{A}\}\,.$ (19)
@@ 5.3 RAY MASK
$g_{\mathcal{A}_{s},\text{RM}}(a)=$ (20)
[…]
where $\varepsilon\ll 1$ ensures numerical stability. The mapping function
$w(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})\colon(0,\lambda_{\mathcal{A}}]^{2}\times\mathbb{R}_{>0}\mapsto(0,1]$ (21)
is chosen such that $\frac{\partial\omega}{\partial\lambda_{a}}>0$, which ensures a convex mapping.
The linear ray mask introduced in [12, Eq. 6] is obtained by setting $w_{\text{lin}}(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})=\frac{\lambda_{a}}{\lambda_{\mathcal{A}}}$.
In addition to the constraints introduced in the problem statement, ray masking requires a star-shaped safe action set to ensure that the safe centre and the line segment from the safe boundary point to the safe centre lie within the set. All convex sets, including zonotopes, are star-shaped. While a safe action set induced by a safe state set is not necessarily convex, it is convex for linearised transition distributions.
Lemma 2.
Let $\mathcal{S}_{s}$ be an explicitly available zonotope and a linearised transition distribution yields $\mathcal{S}_{i+1}(a_{i},s_{i})$ according to Equation˜10. Then, $\mathcal{A}_{s}$ according to Equation˜2 is convex.
We provide a proof in Section˜7.8.
5.3.1 COMPUTATION
The distances to the safe action set can be computed by

### v3

@@ 3.C ZONOTOPES
Translating a zonotope is equivalent to translating the centre. Linearly mapping by $M\in\mathbb{R}^{m\times d}$ yields [46, Eq. 7b]
@@
A support function of a set describes the farthest extent of the set in a given direction. The support function of a zonotope in direction $v\in\mathbb{R}^{d}$ is [62, Lemma 1]
@@
A point $p\in\mathbb{R}^{d}$ is contained in a zonotope if [63, Eq. 6]
@@
Determining the containment of a zonotope in another zonotope is co-NP complete [63], but a sufficient condition for $\mathcal{Z}_{1}\subseteq\mathcal{Z}_{2}$ is [64, Eq. 15]
$1\geq$ $\!\min_{\gamma\in\mathbb{R}^{n_{2}},\Gamma\in\mathbb{R}^{n_{2}\times n_{1}}}$ $\left\lVert\begin{bmatrix}\Gamma&\gamma\end{bmatrix}\right\rVert_{\infty}$ (8a)
@@ 4 PROBLEM STATEMENT
Our work considers constrained Markov decision processes $(\mathcal{S},\mathcal{A},P_{f},r,\mathcal{A}_{s})$ with the following elements:
  • a feasible state set $\mathcal{S}\subseteq\mathbb{R}^{d_{\mathcal{S}}}$,
  • a feasible action set $\mathcal{A}\subseteq\mathbb{R}^{d}$,
@@
We seek a safeguarded, stochastic policy that maximises the expected, discounted return over a finite horizon $N$:
$\pi^{*}(a|s)=\!\operatorname*{argmax}_{\pi(a|s)}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a_{i}\sim\pi(a_{i}|s_{i})\\ s_{i+1}\sim P_{f}\end{subarray}}\sum_{i=0}^{N}\delta^{i}\,r(s_{i},a_{s,i},s_{i+1})$ (9)
with the safe action $a_{s,i}=g(a_{i})$, the continuously differentiable safeguard $g\colon\mathcal{A}\to\mathcal{A}_{s}$, and discount factor $\delta\in(0,1]$.
4.A ENSURING COMPUTATIONAL TRACTABILITY
@@
Safeguarding is computationally cheap when the problem setting provides a specified safe action set. However, safeguarding might be computationally intractable if the safe action set needs to be constructed from a safe state set, as specified in Equation˜2, even if the safe state set is available as a closed, convex set. In these cases, we assume that the next state set can be derived using disciplined convex programming [65]. This makes it possible to replace any constraint on a safe action $a_{s,i}\in\mathcal{A}_{s}$ by the state constraint $\mathcal{S}_{i+1}(a_{s,i},s_{i})\subseteq\mathcal{S}_{s}$.
@@
One method we discuss in particular is obtaining safe state sets via robust control invariant sets [46], which guarantee the existence of an invariance-enforcing controller that can keep all future states within the safe set. This is achieved by enclosing the dynamics at the current state by a linear transition function with a noise zonotope $\mathcal{W}=\left<c_{\mathcal{W}},G_{\mathcal{W}}\right>\subset\mathbb{R}^{d_{\mathcal{S}}}$, such that:
$\mathcal{S}_{i+1}(a_{i},s_{i})=Ma_{i}\oplus\left<c+c_{\mathcal{W}},G_{\mathcal{W}}\right$>\,, (10)
where $c$ is the offset and $M$ the Jacobian of the linearisation. Such an enclosure can, for example, be obtained using reachset-conformant identification [44].
@@ 5 METHOD
Figure˜1 shows the general framework for provably safe, analytic gradient-based reinforcement learning. For any policy output, we apply safeguards that map the unsafe action $a_{i}$ to the safe action $a_{s,i}$. The safe action is executed in the environment, yielding the next state $s_{i+1}$ and reward $r_{i}$. To train the policy, we calculate the gradient of the reward with respect to the policy output as:
$\frac{\partial r_{i}}{\partial a_{i}}=\left(\underbrace{\frac{\partial r_{i}}{\partial a_{s,i}}}_{\text{direct path}}+\underbrace{\frac{\partial r_{i}}{\partial s_{i+1}}\frac{\partial s_{i+1}}{\partial a_{s,i}}}_{\text{indirect path via }s_{i+1}}\right)\frac{\partial a_{s,i}}{\partial a_{i}}\,.$ (11)
@@
Figure 1: The forward (top) pass of the provably safe policy optimisation from analytic gradients describes the integration of the safeguard between the policy and environment. The backward pass (bottom) visualises how we utilise backpropagation to obtain the reward gradient with respect to the policy action. It also highlights the required unrolling of the previous trajectory.
The following subsections detail the safeguard. First, we formulate generally required and desirable properties in the aforementioned differentiable setting. Then, we introduce the two safeguards used in this work: boundary projection (BP) [20] and ray mask (RM) [12]. We structure their introduction by first explaining the general idea of the safeguard, then analysing its properties, and finally presenting our modifications.
5.A REQUIRED AND DESIRED PROPERTIES
A safeguard for our setting must be provably safe as described in Section˜3.B. For backpropagation, it must also guarantee the existence of a Clark generalised derivative [70] everywhere, which implies that the safeguarding is at least of class $C^{0}$.
Beyond the required properties, there are additional desired properties. First, the safeguarding should be of class $C^{1}$ and provide full rank Jacobians $\frac{\partial a_{s}}{\partial a}$ everywhere. A rank-deficient Jacobian can diminish the learning signal by reducing its effective dimensionality, which incurs information loss. Second, the number of interventions by the safeguarding should be minimal throughout training and inference. Minimal interventions also reduce overhead and serve the last desired property of fast computation. In summary, the safeguard must
@@ 5.A REQUIRED AND DESIRED PROPERTIES
    be of class $C^{1}$ and provide full rank Jacobians everywhere, i.e. be a local diffeomorphism,
@@
Subsequently, we present two safeguards that offer different trade-offs between the desired properties. We summarise the properties of the safeguards in Table˜1.
@@
   Specified $\mathcal{A}_{s}$ 1 Quadratic Program 1 Linear Program
   Derived $\mathcal{A}_{s}$ 1 Quadratic Program (1 Conic $\vee$ 1 Quadratic) $\wedge$ 1 Linear Program
5.B BOUNDARY PROJECTION
The boundary projection safeguard, proposed in [20], maps any action to the closest safe action. By definition, it therefore only affects unsafe actions, which are mapped to the nearest boundary point in the safe action set. In Euclidean space, this corresponds to an orthogonal projection to the boundary of the safe action set. We show an exemplary mapping with boundary projection from an unsafe action $a$ to a safe action $a_{s}$ in Figure˜2.
[…]
@@ 5.B BOUNDARY PROJECTION
The safeguard $g_{\text{BP}}(a)$ provides the safe action by solving
@@
5.B.1 PROPERTIES
The optimisation problem is always solvable given Equation˜1, and Equation˜12b ensures satisfaction of Property˜P1. The implicit function theorem [49, Theorem 3.3.1] provides the Jacobian of the solution mapping to satisfy Property˜P2.
The distance between the initial and mapped action decreases smoothly with the distance from the unsafe action to the boundary until it is zero for safe actions. However, the mapping location can change abruptly between unsafe actions on different sides of the edges of the safe set. This leads to a jump in the gradient, such that it is only $C^{1}$ almost everywhere. We employ any element from the Clarke subdifferentiable [70] at non-differentiable points, ensuring our approach remains well-defined. In addition, all actions along the ray starting at a boundary point in the direction of the outward normal are mapped to that boundary point. Formally, any unsafe action $a_{u}$ that can be written as
$\forall t>0\,:a_{u}=a_{s,\partial\mathcal{A}_{s}}+t\cdot v$ (13)
with the safe action on the boundary $a_{s,\partial\mathcal{A}_{s}}\in\partial\mathcal{A}_{s}$ and $v$ any outward normal vector at $a_{s,\partial\mathcal{A}_{s}}$, is mapped by Equation˜12 to $a_{s,\partial\mathcal{A}_{s}}$. Therefore, the safeguard cannot propagate gradients in the mapping direction, such that
$\left(\frac{\partial r}{\partial a_{s}}\frac{\partial a_{s}}{\partial a}\right)v=0\,$, (14)
which is especially problematic for gradients parallel to $v$. In such a case, boundary projection eliminates the gradient, keeping the optimisation stuck indefinitely.
Theorem 1.
Let $\mathcal{A}_{s}$ be the zonotope $\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\right$> with generator matrix $G_{\mathcal{A}_{s}}\in\mathbb{R}^{d\times n}$, such that strict complementary slackness holds for Equation˜12, and $g_{\text{BP}}$ be differentiable. Then the rank of the Jacobian of Equation˜12 is
$\text{rank}\left(\frac{\partial a_{s}}{\partial a}\right)=\begin{cases}d&\text{if }a\in\mathcal{A}_{s}\\ <d&\text{else}\,.\end{cases}$ (15)
We obtain a proof by differentiating through the Karush-Kuhn-Tucker (KKT) conditions, which we provide in Appendix˜B. Consequently, boundary projection does not satisfy Property˜P3.
Boundary projection only intervenes for unsafe actions and therefore adheres to Property˜P4. If the safe action set is specified, Equation˜7 is the containment Equation˜12b. Otherwise, we use the constraint $\mathcal{S}_{i+1}(a_{i},s_{i})\subseteq\mathcal{S}_{s}$, which is tightened by Equation˜8 and convex for a linearised transition function, as Equation˜8 is a linear constraint. Both yield quadratic programs, which compute quickly.
5.B.2 MODIFICATIONS
To regain a gradient in the mapping direction and compensate for the resulting rank-deficient Jacobian, which violates Property˜P3, we augment the policy loss function $l_{r}(a_{s},s)$ with a regularisation term [18, Eq. 16]
@@ 5.B.2 MODIFICATIONS
As a result, the corresponding gradient
$\frac{\partial l}{\partial a}=\frac{\partial l_{r}}{\partial a}+2c_{d}(a_{s}-a)^{T}\left(\frac{\partial a_{s}}{\partial a}-I\right)$ (17)
points along the projection direction $a_{s}-a$. The coefficient $c_{d}$ scales the regularisation to remain small relative to the original loss $l_{r}(a_{s},s)$, yet large enough to produce a meaningful gradient in the mapping direction. In addition to gradient augmentation, the regularisation encourages the policy to favour safe actions from the start, which is desirable as stated in Property˜P4.
5.C RAY MASK
[…]
Figure 3: The ray mask maps actions towards the safe centre $c_{\mathcal{A}_{s}}$ in proportion to the safety domain length $\lambda_{\mathcal{A}_{s}}$, the feasible domain length $\lambda_{\mathcal{A}}$, and the distance from the action to the safe centre $\lambda_{a}$.
The ray mask [12] maps every action radially towards the centre of the safe action set $c_{\mathcal{A}_{s}}$, as shown in Figure˜3. Using the unit vector $d_{a}=\frac{a-c_{\mathcal{A}_{s}}}{\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert_{2}}$, we define the distances from the safe action set centre to the initial action and the boundaries of the safe and feasible action sets as
$\lambda_{a}$ $=\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert_{2}$ (18)
$\lambda_{\mathcal{A}_{s}}$ $=\max\{\lambda\geq 0\mid c_{\mathcal{A}_{s}}+\lambda d_{a}\in\mathcal{A}_{s}\}$ (19)
$\lambda_{\mathcal{A}}$ $=\max\{\lambda\geq 0\mid c_{\mathcal{A}_{s}}+\lambda d_{a}\in\mathcal{A}\}\,.$ (20)
@@ 5.C RAY MASK
$g_{\text{RM}}(a)=\begin{cases}c_{\mathcal{A}_{s}}&\text{if }\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert_{2}<\epsilon\\ c_{\mathcal{A}_{s}}+\omega\lambda_{\mathcal{A}_{s}}d_{a}&\text{else,}\end{cases}$ (21)
where $\varepsilon\ll 1$ ensures numerical stability, and the mapping function, whose arguments were omitted for clarity, $\omega(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})$ is characterised by:
$\omega(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})\colon(0,\lambda_{\mathcal{A}}]^{2}\times\mathbb{R}_{>0}\mapsto(0,1]$ (22)
$\frac{\partial\omega(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})}{\partial\lambda_{a}}>0\,$, (23)
which ensures a safe, convex mapping. The linear ray mask introduced in [12, Eq. 6] is obtained by setting $w_{\text{lin}}(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})=\frac{\lambda_{a}}{\lambda_{\mathcal{A}}}$.
In addition to the constraints introduced in the problem statement in Section˜4, ray masking requires a star-shaped [71, Def. 5.2.9] safe action set to ensure that the safe centre and the line segment from the safe boundary point to the safe centre lie within the set. All convex sets, including zonotopes, are star-shaped. While a derived safe action set is not necessarily convex, it is convex for a linearised transition function.
Theorem 2.
Let $\mathcal{S}_{s}$ be a zonotope and $\mathcal{S}_{i+1}(a_{i},s_{i})$ be the next state set as in Equation˜10. Then, $\mathcal{A}_{s}$ in Equation˜2 is convex.
{IEEEproof}
We start by inserting Equation˜10 into Equation˜2
$\mathcal{A}_{s}(s_{i})=\left\{a_{i}\in\mathcal{A}\mid Ma_{i}\oplus\left<c+c_{\mathcal{W}},G_{\mathcal{W}}\right>\subseteq\mathcal{S}_{s}\right\}\,.$ (24)
The safe action set is convex if and only if the set of all translations $\mathcal{T}=\left\{t\mid t\oplus\left<c+c_{\mathcal{W}},G_{\mathcal{W}}\right>\subseteq\mathcal{S}_{s}\right\}$ is convex. This is the definition of a Minkowski difference, which is convexity preserving [72, Theorem 2.1].
5.C.1 COMPUTATION
The distance to the safe action set $\lambda_{\mathcal{A}_{s}}$ can be computed by